1.5.11 A college foundation raises funds by selling 600 raffle tickets for a new car worth $49,000 at $100 each (a) Find the expected net winnings of a person buying one of the tickets. (b) Find the total profit for the foundation, assuming they had to purchase the car (c) Find the total profit for the foundation, assuming the car was donated (a) The expected net winnings are Round to the nearest hundredth as needed.) Enter your answer in the answer box and then click Check Answer parts remaining Clear All
Answer #1

(a)

\\expected\ net\ earning\ of\ a\ person\\ \\ = probability (winning)*prize+probability(not\ winning)*0\\ \\ = \frac{1}{600}*49000+\frac{599}{600}*0\\ \\=\$81.67\\

(b)

\\profit\ of\ foundations\\ \\ = revenue\ from\ ticket - car\ price \\ \\=600*100-49000\\ \\=\$11000

\\profit\ of\ foundations\\ \\ = revenue\ from\ ticke\\ \\=600*100\\ \\=\$60000

please upvote the answer.